A chemist has two acid solutions.One contains 15 percent acid and the other contains 6 percent acid. How many cubic meters of each solution should be used to obtain 400 cm^3 of a solution that is 9 percent acid?
I need this now...please and thank you

Answer by scott8148(6628)

let x=cm^3 of 15%, so 400-x=cm^3 of 6%

(15%)(x)+(6%)(400-x)=(9%)(400) __ multiply by 100 to clear percentages

15x+2400-6x=3600 __ subtracting 2400 __ 9x=1200 dividing by 9 __ x=400/3

substituting __ 400-(400/3)=800/3